OpenStudy (anonymous):

Find equations of the tangent plane and normal line to the surface z = arctan(y/x) at the point P(1,1,pi/4) Please show work and thank you

OpenStudy (badhi):

First lets find the gradiant of, $$\phi=\tan z -\frac yx$$ $$\nabla\phi=\left \langle \frac y{x^2},-\frac 1x,\sec ^2 z \right\rangle$$ since this is perpendicular to the surface- $$z=\tan^{-1} \frac yx$$ the perpendicular vector to the surface at p=(1,1,pi/4) $$\vec n=\nabla \phi_{(1,1,\frac \pi 4)}=\langle1,-1,2\rangle$$ equation of a plane with a perpendicular vector n, and a point p on the surface is, $$\vec n \cdot(\vec r-\vec p)=0$$ where $$\vec r=\langle x,y,z \rangle$$ equation of the normal line can be taken from, $$t(\vec r-\vec p)=\vec n$$